If 1176 grams of FeS2 is allowed to react with 704 grams of O2 according to the following unbalanced equation, how many grams of Fe2O3 are produced?

FeS2 + O2 → Fe2O3 + SO2
Answer #1

Ans:

Mass of FeS2 taken = 1176 g

Molar mass of FeS2 = 119.98 g/mol

No. of moles of FeS2 taken = 1176/119.98 = 9.80 mol

Mass of O2 taken = 704 g

Molar mass of O2 = 32 g/mol

No. of moles of O2 taken = 704/32 = 22.00 mol

Balanced reaction is: 4 FeS2 + 11 O2 → 2 Fe2O3 + 8 SO2

We can see from the above balanced reaction that for 4 mol of FeS2 we need 11 mol of O2.

So, for 9.80 mol of FeS2, moles of O2 needed = (9.80 x 11)/4 = 26.95 mol. But we have only 22 mol of O2. Meaning, O2 is a limiting reagent.

11 moles of O2 reacts with 4 moles of FeS2. So, 22 moles of O2 reacts with = (22 x 4)/11 = 8 moles of FeS2. Meaning, of the 9.80 moles of FeS2 only 8 moles will react.

We can see from the above balanced reaction that 4 moles of FeS2 produce 2 moles of Fe2O3. So, 8 moles of FeS2 produce 4 moles of Fe2O3.

Molar mass of Fe2O3 = 159.69 g/mol

Mass of 4 moles of Fe2O3 = 159.69 g/mol x 4 mol = 638.76 g

Mass of Fe2O3 produced = 638.76 g